OpenStudy (anonymous):

y'' + 16y = {(2t, 0<=t<=8) , (0, t => 8)} and: y(0) = 0 , y'(0) = 0 find the Laplace transform. the left side is equal to: Y(s^2 + 6) how can i construct the appropriate laplace transform for the right side?

OpenStudy (anonymous):

You can construct the right side out of a modified unit-step (heaviside) functions

OpenStudy (anonymous):

right but how? like u(t)-u(t-8) or something?

OpenStudy (anonymous):

2t*h(t-8) I believe

OpenStudy (anonymous):

does the second inequality not matter?

OpenStudy (anonymous):

2t-2t*h(t-8)

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

You want the function to be 2t until the heaviside "turns on" at 8. At that point the heaviside takes the value that is multiplied to it (2t) and it is subtracted from 2t giving you 0 after 8

OpenStudy (anonymous):

so from 0 to 8, it goes to 2t. from 8 onwards it goes to 0. the heaviside equation jumps from 0 to 1 at 8?

OpenStudy (anonymous):

Yes heaviside would normally take on the value 1 at t=8 onwards. But since were multiplying it by 2t it'll take on that value instead.

OpenStudy (anonymous):

oh and because anything less than 0 is not defined it starts AT 2t

OpenStudy (anonymous):

The heaviside starts at 0 but we have 2t until then by having the heaviside subtract it

OpenStudy (anonymous):

Oh sorry yeah it "starts" at 8, before then is zero

OpenStudy (turingtest):

what was the issue here?

OpenStudy (turingtest):

your transform on the left is wrong as well

OpenStudy (anonymous):

oh yea its Y(s^2 + 16) but my problem is I dont know how he arrived to that answer. Its a little confusing

OpenStudy (anonymous):

to be clearer, i don't know the method to set up the unit step function.

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

@jamej

OpenStudy (turingtest):

you don't understand the left side either?

OpenStudy (turingtest):

http://tutorial.math.lamar.edu/Classes/DE/Laplace_Table.aspx let's use this table according to #36 we get\[s^2Y(s)-sy'(0)-y(0)-Y(s)=s^2Y-Y=Y(s^2-1)\]

OpenStudy (turingtest):

now on the right side what do we know we need to start out with in the equation?

OpenStudy (turingtest):

@babacusazn ^

OpenStudy (anonymous):

the fact that the unit step function is 0 at t<c and 1 at t>c

OpenStudy (anonymous):

well the left side is simple, reduced it is Y(s^2 + 16) = ?

OpenStudy (anonymous):

because the conditions make the s(Y) + y'(0) = 0

OpenStudy (turingtest):

ok, but the right has an inconsistent definition already you have f defined as both 0 and 2t at x=8 is it\[f=2t|t<8\]\[f=0|t\ge8\]?

OpenStudy (anonymous):

exactly oh i made a type 2t, \[0 \le t <8\]

OpenStudy (turingtest):

so that is what I meant above so would you agree then that just looking at \(t<8\) we have \(2t\) on the right?

OpenStudy (turingtest):

\[y''+16y=2t\]covers us up until \(x=8\), agreed?

OpenStudy (anonymous):

correct because it t is not greater than 8, it is not zero

OpenStudy (anonymous):

right y'' + 16' = 2t

OpenStudy (turingtest):

but when it hits \(x=8\) we need an "off" switch we do that by turning "on" the \(inverse\) of \(2t\) at \(x=8\) i.e. we turn on \(-2t\) at \(x=8\)

OpenStudy (turingtest):

x's and t's are all mixed up above :S imagine all t's

OpenStudy (anonymous):

yea no problem

OpenStudy (turingtest):

how can we turn on the inverse of \(2t\) at \(t=8\) using a step function?

OpenStudy (anonymous):

so -2t to turn it off at t = 8 or above that

OpenStudy (turingtest):

right, and how do we write that then?

OpenStudy (anonymous):

u(t-c)?

OpenStudy (anonymous):

c being the point of "inflection" /jump

OpenStudy (turingtest):

c=? here?

OpenStudy (anonymous):

8 :)

OpenStudy (anonymous):

hold on a second, doesn't that mean at c=8, u(t-8) = ehh 1? or

OpenStudy (turingtest):

yes 8

OpenStudy (turingtest):

and so the final expression is...?

OpenStudy (anonymous):

because the unit step function is defined that at 0 t<c and 1 at t>c (or probably \[\ge\])

OpenStudy (turingtest):

befor the laplace I mean

OpenStudy (anonymous):

2t - 2t(u(t-8))

OpenStudy (anonymous):

and from there its a simple look at the table and invert it

OpenStudy (turingtest):

yup :)

OpenStudy (anonymous):

I see, could I post a nother question after I have attempted it? you cleared up the unit step function so clearly thanks!

OpenStudy (turingtest):

I'm goin' to bed bro, it's late for me but ping me when you need help and I'll do my best :) g'night

OpenStudy (anonymous):

thanks alot! :)

OpenStudy (turingtest):

anytime =)